NAOJ GW Elog Logbook 3.2

I upload a PDF on TAMA RF photodetector.

I upload three PDF files on a system of TAMA Faraday Isolator. I do not have information on Faraday crystal.

Members: Sakai-kun, Manuel,
We made a VI that takes a single input, uploads the array buffer, applies a median and an average filter and returns a filtered output. There are controls for the filter array sizes and switches to enable and disable the filters. We put this VI in the loop of the VI that shows the signals, we filtered the X and Y signals and plot the AC and Phase. The plots show the filtered signals in white and the unfiltered in gray.


I contacted to the company (Japan Laser Coorp. as a japanease agency of Newport). The currect legs can be separated into upper and lower parts. Cheapest way to change the height is to change the lower part of the legs. To get a quotation, about one week is needed. An expected delivering time is about three months.

Minutes of Meeting
November 2nd, 2016
Participants: M. Barsuglia, E. Capocasa, R. Flaminio, Y. Guo, M. Tacca
1. Coating at LMA
Discussion on going.
2. Optical scheme
New optical scheme produced by Matteo and Eleonora.
2a. Which polarization shall we use? The OPO and the SHG seems to work with S-pol (TBC with
Matteo Leonardi). Which polarization one is compatible with the EOM that we have?
Yuefan wil check.
2b. SHG lock. It can be done in transmission. GEO did it in transmission. Virgo is doing
it in reflection. If so we can remove the optics to do the lock-in in reflection.
2c. Share of IR power between SHG path and 1064IMC path? Need to check with Matteo Lombardi
what is the minimum power required for the SHG.
2d. Modulation frequency for the SHG/1064MC: GEO was using 75MHz. Virgo is using 80 MHz.
We are planning to use 15 MHz.
2e. IMC's. Mechanical design received form JP Zendri. Finesse: 500? Polarization: S?
Raffaele will consult with Matteo Lombardi. APC team will consult with people at LKB.
2f. Telescopes. Now that the optical lenght are fixed, telescopes can be studied.
Yuefan is studying the telescope for the SHG. Matteo Tacca will pass to Yuefan
the information about three lenses telescopes.
3. Window on vacuum chamber
Flange size is about 15 cm
KAGRA design received from Akutsu. With KAGRA design max window size will be ~9cm.
Matteo Tacca will ask Virgo design for same type of flange.
4. Optical table height. Tatsumi is investigating possible solutions to ahve the beam height
at the same height as the window on the vacuum chamber.
5. Visit to NAOJ of Matteo Tacca. Goal of visit will be the set-up of the green path
for the filter cavity control and the injection into the OPO. Need to define the list of
components to be bought.

Members: Sakai-kun, Manuel
In order to make maps of samples, we made a labview VI that moves the translation stage in 2 directions (namely scan and shift), reads the lockin amp values and put the data in a file.
In attachment the pdf of the front panel and the block diagram.



- Status of assembly
The moving of laser and optics according to Matteo/Eleonora new scheme is started.
We need to check the polarization used/to be used on the bench (is it OK to
work with S-pol?).
Need to check cables behind the lasers. Do they interfere with the beam?
Need to discuss with Matteo and Eleonora.
Yuefan will continue the assembly and fix the lenses required.
- Mode cleaners for 1064 nm and 532 nm
Mechanical design received from Trento. Looking for a company that can produce
them (with the help of Akutsu/ATC).
Curved mirror (1m ROC) may be bought from CVI (Raffaele contacted them)
Trento bought flat mirrors from Opto Sigma.
We can do the same but we need to fix the finesse and the polarization so that
we can fix the coating specifications (to be discussed with Virgo people).
PZT from PI. Raffaele will send the reference to Tatsumi-san to order it.
Viton o-ring to keep the mirror at the good position and allow position control
of the curved mirror. Raffaele will send the reference to Tatsumi-san to order it.
- Height of table and window
Window on vacuum chamber is at 1190 mm. Table is at 900 mm (so beam is at 973 mm).
Three beams need to enter from the window (two squeezed beams at 1064 nm and one
green beam at 532 nm). We would like to avoid periscopes and minimize losses.
So we would like to increase table height by 217 mm so to have the beams at 1190 mm.
Tatsumi-san will investigate what is the solution used in KAGRA.
- Window size
We need to pass three beams through the window.
Yuefan measured the size of the present window.
Size of present window is probably too small.
Larger size is possible. Maybe 9 cm. Is it OK?
The mechanical design used for the KAGRA windows was provided by Autsu-san.
We can start from there. Once the mechancial design will be fixed
we can fix the optic design (diameter, thickness, wedge) and ask
quotation to Sigma Koki.

Sakai-kun found that one of the arms of the translation stage was not moving properly, the step motor was turning but the long black screw wasn't. I found that the grub screw (see the picture) was loose. I checked the other arms and found another loose screw. I guess the company that cleaned and greased it forgot to re-tighten those screws. I tightened them all. Then, by using the Zaber consol, I reset the lockstep of the couples to make them move together again.

1) Glueing Related Things
1-1) Checked the current situation.
One magnet and one stand-off (wire-breaker) should be glued on the BS mirror.
1-2) Got an adhisive of MasterBond EP30-2
Via Hirata-san I got a set of MasterBond EP30-2.
Now it is storaged at TAMA center room.
1-3) Remaining missing part is weight scale.
Takahashi-san will give us. Yuefan, please receive it from Takahashi-san.
1-4) Glueing work will start from 13:30 of 7th November.
2) Suspension Related Things
2-1) Suspension JIG for double pendulm:
Takahashi-san will send back to TAMA from KAGRA site.
An expected delivered date is 9th November (Wed).
2-2) By using dummy mass, Tatsumi will set the hanging wires.
I confirmed that tungsten wires are existing at TAMA.
Expected date for this wotk is 13:30 of 11th November.
2-3) Scheduled date to be install the BS suspension is 16th November.

I did some calculations on the displacement-noise properties of PR3 by using the yaw-noise data measured in April this year from Akutsu-san on the respective OpLev. These calculations should show how the outcome of a length-sensing OpLev-QPD would be influenced by the noise of itself and its surroundings.
It seems by using a lens with f=500 mm we would reach the requirement of measuring at least 1um shift of the mirror. The big feature around 1Hz is due to the Kumamoto earthquake that happend in the period of data-taking.
However, this are just short-term spectra and still it is not quite clear whether such a lens can be installed.
please note that by using a lens with a smaller focal length, the noise spectrum would increase with a factor (900-f)/f, where 900 is the distance between mirror and lens (in mm as f is given in mm too).

Simulations have been done assuming a misplacement of the WAB of 2mm in x and y direction (z is the direction of the main beam) and 1 degree rotation around x and y axis.
The assumptions of shift and rotation are based on the maximum tolerable misplacement of any baffle in KAGRA.
Result: no important impact on the sensitivity of the interferometer has been found.
However, the effect on the power due to the rotation turns out to be 10000 times stronger than due to the shifting!


I ordered the attached parts for SHG housing.
Expected deliver date is 28 October, 2016.

The installation of an "upgraded" version of the OpLevs toward a low-frequency yaw- and length-sensing kind of OpLev has to be done for all main mirrors in KAGRA. However, the PR2 chamber is different from all the other chambers as there is only one viewport available for the (external) OpLevs.
In order to measure the reflected OpLev beam, we either have to install the beam collimeter inside the PR2 chamber so that the reflected beam goes through the only viewport outside, or to install a mirror in the inner wall (proper positioned), so that the beam, coming from outside, is subsequentally reflected directly toward the viewport.
I checked therefore the feasiblity of the second option by using the actual (as I hope) drawings of the PR2 chamber and the payload. It seems that, taken only the drawings, there is no reason why it shouldn't work. I prepared some pictures to see the paths of the OpLev beam (green line) within the chamber. By installing a mirror, it may be possible to guide the reflected beam toward the viewport.
However, there is still an open question how much space the additional structure on the inner wall (small mirror + mount) will take and whether or not it will collide with the main beam passing PR2 chamber toward the BS.

I packed our SHG housing for modification by a company.
Both PZT (mirror) box and heat sink are remained at NAOJ.

I upload a PDF file of
drawings and Parts list for SHG housing

I got a picture from Manuel.